# Private Fine-tuning of Large Language Models with Zeroth-order Optimization

### v1

@@ Abstract
Fine-tuning large pretrained models on private datasets may run the risk of violating privacy. Differential privacy is a framework for mitigating privacy risks by enforcing algorithmic stability. DP-SGD enables training models with private data in a privacy-preserving manner, but raises new obstacles in the form of performance loss and significant engineering challenges. We introduce DP-ZO, a new method for fine-tuning large language models that preserves the privacy of training data by privatizing zeroth-order optimization. A key insight into the design of our method is that the direction of the gradient in SPSA, the zeroth-order algorithm we use, is always random and the only information that depends on private data is the step size, i.e., a scalar. Therefore, we only need to privatize the scalar step size, which is memory-efficient. DP-ZO, which can be instantiated with either Laplace or Gaussian noise, provides a strong privacy-utility trade-off across different tasks, and model sizes, under conservative privacy budgets. One noteworthy result is that DP-ZO exhibits just $1.86\%$ performance degradation due to privacy at $(1,10^{-5})$-DP when fine-tuning OPT-66B on 1000 training samples from SQuAD.
@@ 1 INTRODUCTION
The proliferation of open-source models pretrained on web-scale datasets (Brown et al., 2020; Zhang et al., 2022; Touvron et al., 2023) has created a paradigm shift in privacy preserving machine learning. Differential Privacy (DP) (Dwork et al., 2006) is the gold standard for preserving privacy while training models on private data, but it requires additional data (Tramèr and Boneh, 2021) to prevent a drop in utility (Yu et al., 2021a). Pretrained model checkpoints have emerged as a compelling “free” source of prior information to boost the performance of DP training (Ganesh et al., 2023; Tang et al., 2023a; Panda et al., 2022). By only requiring DP during the fine-tuning phase, a recent line of work (Li et al., 2022b, a; Yu et al., 2021b; He et al., 2023; Bu et al., 2023c) is able to obtain impressive performance with DP-SGD (Abadi et al., 2016). Despite these advancements, DP needs additional engineering effort and may not be available in real-world deployments of fine-tuning APIs (OpenAI, 2023; Anyscale, 2023). We propose a new direction for DP fine-tuning of large pretrained models that achieves strong privacy-utility trade-off and is more resource-efficient, easy to implement, and portable.
In this work, we introduce a new methodology for DP fine-tuning of large pretrained models. Our method uses zeroth-order optimization (ZO) (Spall, 1992). A key insight into this ZO optimization method is the power of two forward passes, each done with the same random perturbation with flipped signs. The difference in losses between the two perturbations tells us which perturbation is better, and we update the model by taking a step in that direction. This approach, which we describe in detail in Section 2.2, allows us to approximate the gradient without backpropagation.
@@
Our differentially private zeroth order (DP-ZO) method privatizes the zeroth-order update, by adding noise to the difference between the losses. This noise is proportional to the sensitivity of this loss difference with respect to changing a single example in the training set, which is controlled by clipping. We limit the $\ell_{p}$ sensitivity by clipping the norm of the difference in scalar losses, between the two random perturbations. By removing the need for per-example gradient clipping (Abadi et al., 2016), our method enables DP training of large models with just a few lines of code.
[…]
@@ 2.1 Differential Privacy
where $D$ and $D^{\prime}$ are neighboring datasets if they differ in a single entry, $\varepsilon$ is the privacy budget, and $\delta$ is the failure probability. When $\delta=0$, we term it as pure $(\varepsilon,0)$-DP or $\varepsilon$-DP for simplicity.
We define a set of existing differential private mechanisms that we will use in our work.
@@
where $z\sim N\left(0,\frac{2\ln(1.25/\delta)\Delta^{2}}{\varepsilon^{2}}\right)$, provides $(\varepsilon,\delta)$-DP.
@@
[…]
@@ 2.2 Zeroth-order Optimization
$\hat{\nabla\mathcal{L}_{b}}(\theta;\mathcal{B})=$ $\frac{\mathcal{L}(\theta+\phi z;\mathcal{B})-\mathcal{L}(\theta-\phi z;\mathcal{B})}{2\phi}z$ (2)
Random perturbations in zeroth-order optimization (ZO) serve as high-variance estimates of the actual gradient, enabling optimization without the need for explicit gradient computations. However, these perturbations themselves carry a privacy risk. The characteristics of the perturbations can be inferred from the gradient updates, effectively leaking information about the data. Moreover, the magnitude of the perturbations can be estimated with high precision in high dimensions due to the central limit theorem. Therefore, incorporating differential privacy into ZO is essential to safeguard against these vulnerabilities.
@@ 3 OUR METHOD: DP-ZO
We introduce our framework for differentially private zeroth order optimization (DP-ZO) by integrating DP into Definition 2.4. In our framework, the information obtained from training data can be represented as a scalar. This scalar has a bounded sensitivity (when applying clipping) and can be privatized by adding noise. We expect this noise to be less detrimental to the utility, compared to the noise added in DP-SGD, as we are adding noise to a single dimension. In other words, we would expect the gap between non-private and private utility to be smaller than that of DP-SGD. However, it is possible that for some tasks the non-private performance of zeroth-order optimization is poor.
@@
12:     $s=\frac{\sum_{i\in\mathcal{B}}l_{i}+\mathcal{Z}(\mathrm{C},\sigma)}{B\cdot 2\phi}$
13:     $\theta=\theta-\eta s\vec{z}$
DP-ZO.
@@
We explain the steps of our algorithm while emphasizing the key differences from Definition 2.4 required to guarantee $(\varepsilon,\delta)$-DP. We first sample a batch from the dataset with Poisson sampling (Balle et al., 2018) which allows us to use privacy amplification by subsampling. For each model parameter $\theta_{i}$ we want to update, we independently sample a perturbation $z_{i}$ from a standard Gaussian distribution and scale it by a predetermined constant $\phi$; we denote the full perturbation vector as $\phi\vec{z}$. Now we compute an approximation of the gradient by projecting it onto the random perturbation $\vec{z}$. That is, for a training sample $x_{i}$ we compute the difference in scalar losses between ${\theta+\phi\vec{z},\theta-\phi\vec{z}}$. Intuitively, this scalar tells us how much better one random step is than the other. We clip this scalar to limit the sensitivity (described in detail in the next paragraph). We add noise to the aggregation over samples in our training batch (described in detail in the subsequent paragraph). Finally, we take a step in the direction of $\vec{z}$ by scaling our private step size by the expected batch size, perturbation constant $\phi$, and the learning rate $\eta$.
DP-ZO enables new mechanisms by privatizing the difference in losses between perturbations.
@@
DP-ZO proposes an update direction determined by a $d$-dimensional random vector, generated from a public PRNG, making the direction itself public information. The only private aspect is the step size, that is influenced by the difference in losses between perturbations with opposite signs. To privatize this step size, we add noise proportional to the sensitivity of the step size. We bound the sensitivity of the step size by clipping the per-example step sizes to a specific range $[-C,C]$, so the sensitivity under add-remove DP is $C$.
@@
Given a private scalar with bounded sensitivity, we can apply the classical Gaussian mechanism to release a privatized scalar with $(\varepsilon,\delta)$-DP. The Gaussian mechanism is widely studied in privacy-preserving machine learning techniques like DP-SGD, in part because the best accounting techniques for the Gaussian mechanism (Dong et al., 2019; Gopi et al., 2021) are tight. However, the Gaussian mechanism can only provide $(\varepsilon,\delta)$-DP. To avoid such a failure case in extremely sensitive applications, researchers often recommend using cryptographically small values of $\delta$ (Vadhan, 2017). Unfortunately, due to limitations of accounting methods, we currently cannot calculate the tight privacy of composition of sub-sampled Gaussian mechanism for values of $\delta$ smaller than $10^{-10}$. Alternatively, we can resort to mechanisms that can obtain pure $\varepsilon$-DP. These mechanisms, such as Laplace mechanism, come with a guarantee that the mechanism will never fail catastrophically. However, due to large tails of the Laplace mechanism, it has never been a contender for high dimensional optimization.
Although it is possible to obtain pure DP with DP-SGD by adding Laplace noise scaled to the $\ell_{1}$ sensitivity of the gradient, this is challenging for large models because the $\ell_{1}$ sensitivity can be $\sqrt{d}$ times larger than the $\ell_{2}$ sensitivity (and often is), especially in the modern large machine learning models that consist of billions of parameters. In contrast, DP-ZO only requires to privatize the loss. The one-dimensional private estimation of the step size is amenable to the Laplace mechanism, because the $\ell_{p}$ norms are equivalent.
@@
Specifically, DP-ZO with the Laplace mechanism is the first method to achieve a reasonable privacy-utility trade-off under pure $\varepsilon$-DP for private fine-tuning of LLMs. While this work primarily explores these two mechanisms, the DP-ZO framework is flexible enough to be extended to other differential privacy mechanisms, broadening its applicability.
Privacy Analysis.
@@
As we consider multiple accounting methods with multiple previously proposed mechanisms, we give the overview of the analysis below and defer the full privacy analysis to Appendix A.
@@
Proof Overview. We analyze the privacy of models trained iteratively with DP-ZO by composing the per-iteration privacy loss (Kairouz et al., 2015). At each step, line 11 upper bounds the $\ell_{1}$ (and therefore $\ell_{p}\forall p\geq 1$) sensitivity of the difference in losses $l_{i}$ by $C$. Line 12 adds noise based on DP mechanisms such that each step in Algorithm 1 satisfies $(\varepsilon,\delta)$-DP with some privacy parameters. Therefore Algorithm 1 is $(\varepsilon,\delta)$-DP with some privacy parameters that are calculated via some mechanism-dependent composition theorem. The specific $\varepsilon,\delta$ depends on the choice of mechanism. Most of the privacy analyses in Appendix A are based on the numerical composition of private random variable for the corresponding privacy curve (Gopi et al., 2021) except for the pure $\varepsilon$-DP analysis. The analysis for pure $\varepsilon$-DP by the Laplace mechanism is based on the basic composition (Dwork and Roth, 2014). Given a number of iterations $T$ and the use of the Laplace mechanism $Laplace(0,\sigma)$, Algorithm 1 is $\varepsilon$-DP with $\varepsilon=T\cdot\log(1+p\cdot(e^{1/\sigma}-1))$.
@@
Remark. We can get a tighter privacy composition of $\varepsilon$ by relaxing the Laplace mechanism’s $\varepsilon$-DP guarantee to $(\varepsilon,\delta)$-DP guarantee. We can compute the composition of privacy loss for Laplace mechanism by the private random variable for privacy curve of a $\varepsilon$-DP algorithm in Gopi et al. (2021). We also provide a more fine grained privacy analysis by using the dominating pair for the subsampled Laplace mechanism in the accounting method of Gopi et al. (2021) (See Appendix A).

### v2

@@ Abstract
Differentially private stochastic gradient descent (DP-SGD) allows models to be trained in a privacy-preserving manner, but has proven difficult to scale to the era of foundation models. We introduce DP-ZO, a private fine-tuning framework for large language models by privatizing zeroth order optimization methods. A key insight into the design of our method is that the direction of the gradient in the zeroth-order optimization we use is random and the only information from training data is the step size, i.e., a scalar. Therefore, we only need to privatize the scalar step size, which is memory-efficient. DP-ZO provides a strong privacy-utility trade-off across different tasks, and model sizes that are comparable to DP-SGD in $(\varepsilon,\delta)$-DP. Notably, DP-ZO possesses significant advantages over DP-SGD in memory efficiency, and obtains higher utility in $\varepsilon$-DP when using the Laplace mechanism.
@@ 1 INTRODUCTION
Figure 1: Visualization of DP-ZO. The only information from private data is a scalar and we only need to add noise to this scalar. This scalar privatization enjoys the benefits of flexibility with DP mechanisms, ease of implementation, and reduced computation.
@@
The proliferation of open-source models pretrained on web-scale datasets (Brown et al., 2020; Zhang et al., 2022; Touvron et al., 2023) has created a paradigm shift in privacy preserving machine learning. Differential Privacy (DP) (Dwork et al., 2006) is the gold standard for preserving privacy while training models on private data, but it requires additional data (Tramèr and Boneh, 2021) to prevent a drop in utility (Yu et al., 2021a). Pretrained model checkpoints have emerged as a compelling “free” source of prior information to boost the performance of DP training (Ganesh et al., 2023; Tang et al., 2023a; Panda et al., 2024a). By only requiring DP during the fine-tuning phase, a recent line of work (Li et al., 2022b, a; Yu et al., 2022; He et al., 2023; Bu et al., 2023d) is able to obtain impressive performance with DP-SGD (Abadi et al., 2016). Despite these advancements, DP-SGD causes additional memory cost and needs additional engineering effort, especially for large models across devices. We propose a new direction for DP fine-tuning of large pretrained models that achieves strong privacy-utility trade-off and is more resource-efficient, easy to implement, and portable.
@@
In this work, we introduce a new methodology DP-ZO for DP fine-tuning of large pretrained models. Our method uses zeroth-order optimization (ZO) (Spall, 1992). Our key insight is the synergy between differentially private fine-tuning and zeroth-order optimization. ZO provides the gradient estimates and the only information from private data in ZO is a scalar. We only need to privatize the scalar update by adding noise to it. Specifically, the scalar is the differences between losses from models with the same random perturbation but flipped signs. DP-ZO privatizes the zeroth-order update, by adding noise to the difference between the losses (visualized in Figure 1). This noise is proportional to the sensitivity of this loss difference with respect to changing a single example in the training set, which is controlled by clipping. We limit the $\ell_{p}$ sensitivity by clipping the norm of the difference in scalar losses, between the two random perturbations. Therefore, DP-ZO is flexible for both $\varepsilon$-DP and $(\varepsilon,\delta)$-DP. By removing the need for per-example gradient clipping (Abadi et al., 2016), DP-ZO enables DP training of language models with just a few lines of code and without the need for backpropagation.
Figure 2: DP-ZO provides a strong privacy-utility trade-off across different tasks under conservative privacy budgets.
Figure 3: DP-ZO achieves comparable performance as DP-SGD with same model size and scales seamlessly to large models like 30B/66B, that are challenging for DP-SGD.
Figure 4: DP-ZO achieves non-trivial performance for $\varepsilon$-DP. In contrast, DP-SGD (laplace) suffers to improve upon $\varepsilon=0$ (zero-shot) due to high variance.
DP-ZO provides a strong privacy-utility trade-off across different tasks, model sizes, dataset sizes, and DP mechanisms under conservative privacy budgets. DP-ZO only slightly degrades the performance compared to the non-private baseline (Figure 4). DP-ZO achieves comparable performance as DP-SGD within the same model size from 1.3B to 13B (Figure 4). DP-ZO scales seamlessly to large models without additional engineering, while DP-SGD requires much more memory and effort to implement per-example gradient clipping across GPUs (within a reasonable research computation limit, DP-SGD results on OPT-30B/66B are not available and omitted in Figure 4). As the model size increases to OPT-66B, the performance of DP-ZO increases and the utility gap between DP-ZO and the non-private baseline also decreases (Figure 4). Because our method only privatizes a scalar, it is compatible with multiple DP mechanisms. Specifically, DP-ZO is the first method to provide pure $\varepsilon$-DP with nontrivial utility (73.52 for SQuAD at $\varepsilon=4$) for large models by using the Laplace mechanism (Figure 4).
Besides, we provide the empirical privacy analysis of ZO and DP-ZO. While ZO itself incurs less empirical privacy loss than SGD, such empirical privacy analysis is still much higher than random guess. DP-ZO can reduce such privacy attack close to random guess. We also show the computation efficiency of DP-ZO over DP-SGD, even when applying gradient checkpointing and half-precision to both methods.
@@ 2.1 Differential Privacy
When $\delta=0$, we term it as pure $(\varepsilon,0)$-DP or $\varepsilon$-DP for simplicity.
We define a set of existing DP mechanisms that we will use in our work.
@@
where $z\sim\mathcal{N}\left(0,\frac{2\ln(1.25/\delta)\Delta^{2}}{\varepsilon^{2}}\right)$, provides $(\varepsilon,\delta)$-DP.
@@
where $z\sim\text{Laplace}\left(0,\frac{\Delta}{\varepsilon}\right)$, provides $(\varepsilon,0)$-DP.
@@ 2.2 Zeroth-order Optimization
^$\hat{\nabla}\mathcal{L}_{b}(\theta;\mathcal{B})=$ $\frac{\mathcal{L}(\theta+\phi z;\mathcal{B})-\mathcal{L}(\theta-\phi z;\mathcal{B})}{2\phi}z$ (2)
Random perturbations in zeroth-order optimization (ZO) serve as high-variance estimates of the actual gradient, enabling optimization without the need for explicit gradient computations. However, these perturbations themselves carry a privacy risk. The characteristics of the perturbations can be inferred from the gradient updates, effectively leaking information about the data. Therefore, incorporating differential privacy into ZO is essential to safeguard against these vulnerabilities.
@@ 3 OUR METHOD: DP-ZO
We introduce our framework for differentially private zeroth order optimization (DP-ZO) by integrating DP into Definition 2.4. In our framework, the information obtained from training data can be represented as a scalar. This scalar has a bounded sensitivity (when applying clipping) and can be privatized by adding noise. If we compare the noise added in DP-ZO to a single dimension to the noise added in DP-SGD to the entire gradient, we expect the univariate noise to be less detrimental to the utility (due to the curse dimensionality in differential privacy Dwork and Roth (2014)). In other words, we would expect the gap between non-private and private utility to be smaller than that of DP-SGD. However, it is possible that for some tasks the non-private performance of zeroth-order optimization is poor (see Section 4.3).
@@
12:     end for
13:     $s=\frac{\sum_{i\in\mathcal{B}}l_{i}+\mathcal{Z}(\mathrm{C},\sigma)}{B\cdot 2\phi}$
14:     $\theta=\theta-\eta s\vec{z}$
15:end for
@@
DP-ZO. We explain the steps of our algorithm while emphasizing the key differences from Definition 2.4 required to guarantee $(\varepsilon,\delta)$-DP. We first sample a batch from the dataset with Poisson sampling (Balle et al., 2018) which allows us to use privacy amplification by subsampling. For each model parameter $\theta_{i}$ we want to update, we independently sample a perturbation $z_{i}$ from a standard Gaussian distribution and scale it by a predetermined constant $\phi$; we denote the full perturbation vector as $\phi\vec{z}$. Now we compute an approximation of the gradient by projecting it onto the random perturbation $\vec{z}$. That is, for a training sample $x_{i}$ we compute the difference in scalar losses between ${\theta+\phi\vec{z},\theta-\phi\vec{z}}$. Intuitively, this scalar tells us how much better one random step is than the other. We clip this scalar to limit the sensitivity . We add noise to the aggregation over samples in our training batch (described in detail in the subsequent paragraph). Finally, we take a step in the direction of $\vec{z}$ by scaling our private step size by the expected batch size, perturbation constant $\phi$, and the learning rate $\eta$.
@@
DP-ZO enables new mechanisms by privatizing the difference in losses between perturbations. DP-ZO proposes an update direction determined by a $d$-dimensional random vector (sampled from standard Gaussian distribution) independent of the private training data. The only private aspect is the step size, that is influenced by the difference in losses between perturbations with opposite signs. To privatize this step size, we add noise proportional to the sensitivity of the step size. We bound the sensitivity of the step size by clipping the per-example step sizes to a specific range $[-C,C]$, so the sensitivity under add-remove DP is $C$.
@@
Given a private scalar with bounded sensitivity, we can apply the classical Gaussian mechanism to release a privatized scalar with $(\varepsilon,\delta)$-DP. The Gaussian mechanism is widely studied in privacy-preserving machine learning techniques like DP-SGD, in part because the best accounting techniques for the Gaussian mechanism (Dong et al., 2019; Gopi et al., 2021) are tight. However, the Gaussian mechanism can only provide $(\varepsilon,\delta)$-DP and researchers often recommend using cryptographically small values of $\delta$ (Vadhan, 2017). Unfortunately, due to limitations of accounting methods, we currently cannot calculate the tight privacy of composition of sub-sampled Gaussian mechanism for values of $\delta$ smaller than $10^{-10}$. Alternatively, we can resort to mechanisms that can obtain pure $\varepsilon$-DP. These mechanisms, such as Laplace mechanism, come with a guarantee that the mechanism will never fail catastrophically. However, due to large tails of the Laplace mechanism, it has never been a contender for high dimensional optimization.
@@
Although it is possible to obtain pure DP with DP-SGD by adding Laplace noise scaled to the $\ell_{1}$ sensitivity of the gradient, this is challenging for large models because the $\ell_{1}$ sensitivity can be $\sqrt{d}$ times larger than the $\ell_{2}$ sensitivity (and often is; see Section 4.3), especially for billion-parameter LLMs. In contrast, DP-ZO only requires privatizing the loss. The one-dimensional private estimation of the step size is amenable to the Laplace mechanism, because the $\ell_{p}$ norms are equivalent. Specifically, DP-ZO with the Laplace mechanism is the first method to achieve a reasonable privacy-utility trade-off under pure $\epsilon$-DP for private fine-tuning of LLMs. While this work primarily explores these two mechanisms, the DP-ZO framework is flexible enough to be extended to other differential privacy mechanisms, broadening its applicability.
@@
Privacy analysis. As we consider multiple accounting methods with multiple previously proposed mechanisms, we give the overview of the analysis below and defer the full privacy analysis to Appendix A.
@@
Proof Overview. We analyze the privacy of models trained iteratively with DP-ZO by composing the per-iteration privacy loss (Kairouz et al., 2015). At each step, line 11 upper bounds the $\ell_{1}$ (and therefore $\ell_{p}\forall p\geq 1$) sensitivity of the difference in losses $l_{i}$ by $C$. Line 13 adds noise based on DP mechanisms such that each step in Algorithm 1 satisfies $(\varepsilon,\delta)$-DP with some privacy parameters. Therefore Algorithm 1 is $(\varepsilon,\delta)$-DP with some privacy parameters that are calculated via some mechanism-dependent composition theorem. The specific $\varepsilon,\delta$ depends on the choice of mechanism. Most of the privacy analyses in Appendix A are based on the numerical composition of private random variable for the corresponding privacy curve (Gopi et al., 2021) except for the pure $\varepsilon$-DP analysis. The analysis for pure $\varepsilon$-DP by the Laplace mechanism is based on the basic composition (Dwork and Roth, 2014). Given a number of iterations $T$ and the use of the Laplace mechanism $Laplace(0,\sigma)$, Algorithm 1 is $\varepsilon$-DP with $\varepsilon=T\cdot\log(1+p\cdot(e^{1/\sigma}-1))$.
@@
Remark. We can get a tighter composition of $\varepsilon$ by relaxing Laplace mechanism’s $\varepsilon$-DP to $(\varepsilon,\delta)$-DP. Note that since we are dealing with scalar values, our mechanism in each iteration will be a one dimensional Laplace mechanism. Therefore we can compute the dominating pair for a single dimensional Laplace mechanism based on Zhu et al. (2022); Wang et al. (2023), that is tighter than directly using the private random variable for privacy curve of a $\varepsilon$-DP algorithm in Gopi et al. (2021). We detail the full privacy analysis in Appendix A.
Proposition 3.2.
DP-ZO attains a convergence rate $\mathcal{O}(\sqrt{r}/\varepsilon n)$, where $r$ is the effective rank of the problem.
Malladi et al. (2023) proves the convergence rate of fine-tuning of language models with zeroth-order optimization is proportional to $r$ instead of the model dimension. A concurrent work (Zhang et al., 2024a) independently proposes DP-ZO and provides the convergence analysis for DP-ZO with Gaussian mechanism, that is independent of the model dimension in private training (Song et al., 2021; Li et al., 2022a). We discuss our work and Zhang et al. (2024a) in Section 6.
4 EVALUATION
We first overview our experimental setup in Section 4.1 and then evaluate the performance of DP-ZO in Section 4.2. We find that DP-ZO provides a competitive privacy-utility trade-off for conservative privacy budgets across multiple datasets, model architectures and can scale to large models under conservative privacy budgets. We also compare DP-ZO to DP-SGD in Section 2 and show that DP-ZO achieves comparable performance to DP-SGD for the same model size. Furthermore, we show that DP-ZO achieves a non-trivial privacy-utility trade-off under pure $\varepsilon$-DP under a conservative privacy budget like $\varepsilon=4$ on large language models. In Section 4.3 we first ablates DP-ZO across different model architectures. We then measure the empirical privacy loss and computation efficiency of DP-ZO. We also characterize DP-ZO under different few-shot settings and different noise mechanisms for $(\varepsilon,\delta)$-DP.
4.1 Experimental Setup
We report the metric of interest (F1 score or accuracy) and standard deviation averaged across 5 independent runs with different random seeds. We detail the full hyperparameter searches in Appendix D.
Datasets. We mainly consider three different benchmark NLP tasks: SQuAD (Rajpurkar et al., 2016) and DROP (Dua et al., 2019) for text generation, and SST2 (Socher et al., 2013) for text classification. Although all these datasets have very different dataset sizes, we consider the few-shot setting for all these datasets where we sample $1000$ points for each dataset. Fine-tuning LLMs with $O(n=1000)$ samples is a standard setting in the NLP community (Gao et al., 2021; Malladi et al., 2023) because we are generally interested in the few-shot abilities of LLMs (Brown et al., 2020). This represents a departure from prior works that privately finetune LLMs; Yu et al. (2022); Li et al. (2022b); Yu et al. (2021b) use the entire training dataset of SST2 that has about 65,000 examples. It is well known that the privacy-utility tradeoff improves greatly with more data (Tramèr and Boneh, 2021). It is straightforward to see that our setting with datasets of the size $n=1000$ with $\delta=10^{-5}$ is simultaneously more challenging and more aligned with real-world usecases than previous works in DP finetuning of LLMs. Despite the increased difficulty of our few-shot setting as compared to prior work, our results validate that DP-ZO realizes a strong privacy-utility trade-off. We also ablates the training sample size from the few-shot to the full training set by conducting experiments on the QNLI (Wang et al., 2019) dataset to be consistent with previous works (Li et al., 2022b; Yu et al., 2022) for a fair comparison.
Models. We present our main results (Table 1) using a pretrained OPT-13B (Zhang et al., 2022) model that is finetuned with LoRA (Hu et al., 2022); that is, we update $<1\%$ of the total parameters. We include a range of ablation studies, including varying the model size among the OPT series, model architectures including Mistral-7B-v1 (Jiang et al., 2023) and amount of parameters to be updated, after we present the main results. We also include one experiments for QNLI on Roberta-base (Liu et al., 2019) to be consistent with previous works (Li et al., 2022b; Yu et al., 2022) for a fair comparison.
Privacy budgets. We consider various privacy levels with $\varepsilon=[0.5,1,4]$ and fix $\delta=10^{-5}$ for $(\varepsilon,\delta)$-DP. We include the zero-shot $\varepsilon=0$ baseline that does not incur any privacy loss because we evaluate the pretrained model directly without finetuning on private data. We also include the non-private $\varepsilon=\infty$ baseline that is trained without any DP guarantee. That is, we iterate over the shuffled dataset instead of doing Poisson sampling (replacing line 4), do not clip the step size (skipping line 11) and set $\sigma=0$ (in line 13). We make these modifications because Poisson sampling and clipping are known to degrade performance, and we want to compare to the strongest possible nonprivate baseline.
4.2 Main Results
DP-ZO provides a strong privacy-utility trade-off for conservative privacy budgets. As shown in Table 1, across all three tasks and all $\varepsilon$s, DP-ZO significantly improves upon the $\varepsilon=0$ baseline, and only slightly degrades the performance compared to the non-private baseline. For SQuAD, even at $\varepsilon=0.5$, DP-ZO can still achieve $80.10\%$, that significantly outperforms $\varepsilon=0$ baseline ($46.23\%$). The gap between $\varepsilon=0.5$ and $\varepsilon=\infty$ is about $6.75\%$. By increasing $\varepsilon$ from $0.5$ to $4$, this gap can be further reduced to $3\%$. For DROP and SST2, DP-ZO (Gaussian) achieves comparable performance as the non-private baseline at $\varepsilon=4$.